OpenStudy (anonymous):

1,296^(2n)=216^(n+10) help please

Parth (parthkohli):

= (6^4)^2n = (6^3)^n+10 = 6^8n = 6 ^ 3n + 30 = 3n + 30 = 8n = 5n = 30 => n = 30/5 n = 6

OpenStudy (anonymous):

parth i think u r correct!!

OpenStudy (tommo_lcfc):

Yes @ParthKohli is correct, I did it the long way and just cancelled the 10/2n to 10/n rather than 5/n.

OpenStudy (anonymous):

\[(6^4)^2n = (6^3)^n+10\] \[ 6^8n = 6 ^ 3n + 30\] \[3n + 30 = 8n\] \[\rightarrow n = 30/5\] \[n = 6\]

Parth (parthkohli):

@Rohangrr check your R.H.S in the first step, the +10 is in the exponent. It's not added to 6^3n

OpenStudy (anonymous):

@ParthKohli its whole exponential!